Two cars are traveling from the same place and in the same direction. The first car started from rest and speeds up with a constant acceleration of 0.3 m/s^2. The second car leaves 15 seconds after the first car and speeds up with a constant acceleration of 0.5 m/s^2. When will the second car overtake the first car?
1
Expert's answer
2020-10-07T10:12:41-0400

"s_0=a_1t^2/2=0.3\cdot15^2/2=33.75(m)"

"v_0=a_1t=0.3\cdot15=4.5(m/s)"


"s_1=s_0+v_0t+a_1t^2\/2"

"s_2=a_2t^2\/2"


"s_1=s_2\to s_0+v_0t+a_1t^2/2=a_2t^2/2"


"33.75+4.5t+0.3t^2\/2=0.5t^2\/2"


"0.1t^2-4.5t-33.75=0" "\to t=51.5(s)"